OpenStudy (anonymous):

Consider the line that passes through the points (4, -3) and (2, -1). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y - 3 = -1(x + 4) y - 1 = -1(x + 2) Choose one answer. a. Only equation #1 represents the line that passes through the two given points. b. Only equation #2 represents the line that passes through the two given points. c. Both equation #1 and equation #2 represent the line that passes through the two given points. d. Neither equation #1 nor equation #2 represents the line that passes

OpenStudy (anonymous):

Equation 2

OpenStudy (anonymous):

Wait a minute.

OpenStudy (anonymous):

Answer is D.

OpenStudy (anonymous):

Equation 1 and Equation 2 are both the same. Neither of them cross through both lines.

OpenStudy (anonymous):

That's what i figure thanks

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) slope = (-1 - (-3) / (2 - 4) slope = (-1 + 3) / -2 slope = 2/-2 = -1 y - y1 = m(x - x1) y - (-3) = -1(x - 4) y + 3 = -1(x - 4) y - y1 = m(x - x1) y - 2 = -1(x - 1) I agree...it is D